### v1

@@ Abstract
We study the emerging collective states in a simple mechanical model of a dense group of self-propelled polar disks confined within a circular arena. Each disk has its center of rotation located a distance $R$ behind its centroid, so its central linear contact forces can produce torques leading to mutual alignment. Each disk is thus affected by this mutual alignment to its neighbors’ orientations and by self-alignment (resulting from disk-substrate interactions) towards the sum of forces over it. The effect of both alignment mechanisms produces a variety of collective states combining high-frequency localized circular oscillations with low-frequency milling around the center of the arena, in fluid or solid regimes. We consider cases with small/large $R$ values, isotropic/anisotropic disk-substrate damping, smooth/rough arena boundaries, different densities, and multiple systems sizes, showing that the identified emergent collective states are robust, generic, and potentially observable in real-world natural or artificial systems.
keywords:
active agents, anisotropy, self-alignment, collective states, vortex, milling
Introduction
@@
A natural way to combine self-alignment and mutual alignment effects in a mechanical model is to consider off-centered attraction-repulsion forces that are not radial with respect to the center of rotation. The inter-agent forces then introduce torques that can result in mutual alignment, while simultaneously producing self-alignment. This approach was implemented in a minimal model of active solid introduced in [54], where a system of self-propelled agents connected by springs attached to level arms was shown to display two types of order-disorder transition, one driven by the focusing of the self-propulsion energy into low elastic modes (produced by self-alignment) and the other by the decentralized heading consensus (resulting from mutual alignment). A similar approach was carried out to show that both types of self-organizing dynamics are also present in a minimal model describing densely packed self-propelled disks with linear repulsion only, each with an off-centered center of rotation, behind its geometrical center [55].
@@
Our work shows that dense systems of self-propelled agents with off-centered rotational dynamics can produce, both, self-alignment and mutual alignment processes, demonstrating the coexistence of two forms of self-organization in active systems. We expect experimental systems to display both types of self-organized dynamics in realistic setups, thus resulting in rich combinations of emergent states.
Results
Dynamics of active polar disks with off-centered rotation
We describe a dense system of active elastic agents confined in a circular arena, to carry out simulations as detailed in the Methods section. Each agent is a soft disk $i$ with position $\vec{r}_{i}$, defined by its center of rotation, that self-propels with preferred speed $v_{0}$ along the heading direction indicated by its orientation unit vector $\hat{n}_{i}=(\cos(\theta_{i}),\sin(\theta_{i}))$, as shown in Fig. 1a. All disks have the same diameter $l_{0}$ and a repulsive force appears between disks $i$ and $j$ when they overlap, given by $\vec{f}_{ij}=(k/l_{0})(|\vec{l}_{ij}|-l_{0})\vec{l}_{ij}/|\vec{l}_{ij}|$ if $|\vec{l}_{ij}|<l_{0}$, and by $\vec{f}_{ij}=0$ otherwise, where $\vec{l}_{ij}=\vec{l}_{j}-\vec{l}_{i}$ with $\vec{l}_{i}=\vec{r}_{i}+R\hat{n}_{i}$. The total repulsive forces exerted by other agents on disk $i$ will thus be $\vec{F}_{i}={\textstyle\sum_{j\epsilon S_{i}}}\vec{f}_{ij}$, where $S_{i}$ includes the indexes of all the neighbors of disk $i$.
Figure 1: Active polar disks with linear elastic repulsion model and arena. a) Schematic diagram of two interacting self-propelled disks, $i$ and $j$, oriented along $\hat{n}_{i}$ and $\hat{n}_{j}$, respectively. Each disk has diameter $l_{0}$ and its center of rotation is labeled by a red point, located a distance $R$ behind its geometrical center. Linear repulsive forces, proportional to $|\vec{l}_{ij}|-l_{0}$, act on each centroid, resulting in forces and torques on $\vec{r}_{i}$ and $\vec{r}_{j}$. b) Snapshot of a relatively ordered densely packed group of active polar disks in circular confinement, inside a smooth circular barrier introduced as a radial linear outer potential. c) Snapshot of a relatively disordered densely packed group of active polar disks confined by a ring of fixed disks with the same repulsion potential as the active disks, representing a rough boundary. Each arrow corresponds to an active disk and its self-propulsion direction, colored by angle.
The combined effect of the interaction forces and self-propulsion leads to the following overdamped equation of motion for the position of each agent [55, 4]
@@ Results
The disk-substrate interaction can lead to self-alignment, where the heading tends to align towards the sum of total forces on the disk. While self-aligning torques may originate from various mechanisms, as discussed in [4], we consider here that they have the form $(\hat{n}_{i}\times\vec{F}_{i})\times\hat{n}_{i}=(\hat{n}_{i}\times\vec{F}_{i})\cdot\hat{z}=(\vec{F}_{i}\cdot\hat{n}_{i}^{\perp})\hat{n}_{i}^{\perp}=(\mathbb{I}-\hat{n}_{i}\hat{n}^{T}_{i})\vec{F}_{i}$, where $\hat{z}$ is the unit vector normal to the plane and $\beta$ controls the self-aligning strength. Note that these torques will depend on the details of the disk-substrate interaction. On the other hand, the position and center of rotation $\vec{r}_{i}$ of each disk $i$ is located a distance $R$ behind its geometrical center (with respect to its heading direction), at $\vec{l}_{i}=\vec{r}_{i}+R\hat{n}_{i}$, which introduces mutual alignment due to the torques on the arm of length $R$ between the geometrical center (where the central forces are acting) and the center of rotation. By combining these torques with the effects of noise, we find a dynamical equation for the heading direction $\hat{n}_{i}$, given by
@@
where $\eta_{i}$ is a random variable that introduces Gaussian white noise with zero mean and variance $\left\langle\eta_{i}(t)\eta_{j}(t^{\prime})\right\rangle=\delta_{ij}\delta(t-t^{\prime})$.
We will study the collective regimes that result from integrating Eqs. (1) and (2) for a dense group of disks confined in a circular arena under different conditions, as shown in the examples in Figs. 1b and 1c.
@@
Overview of emergent states
We will start by providing an overview of the different regimes found for agents with isotropic or anisotropic damping and small or large $R$, confined by a smooth or rough boundary. All simulations were carried out as detailed in the Methods section.
We characterize the emerging collective states using two order parameters. The first one the usual Polarization ($P$). It shows the degree of alignment and is defined by
$P=\frac{1}{N}\langle\|\sum_{i=1}^{N}{\hat{n}_{i}}\|\rangle_{t}~{},$ (3)
where $\left\langle\cdot\right\rangle_{t}$ denotes the average over time, after reaching the steady state. The second one measures the degree of milling in the system, and is defined as the milling order parameter
Here, $\hat{r}_{ig}=\left(\vec{r}_{i}-\vec{r}_{g}\right)/\left\|\vec{r}_{i}-\vec{r}_{g}\right\|$ points to the position of agent $i$ with respect to the center of the group $\vec{r}_{g}$.
Figure 2 shows the collective states observed in our simulations of 400 disks with diameter $l_{0}=1$, self-propulsion speed $v_{0}=0.002$, alignment strength $\beta=1.2$, low noise $D_{\theta}=0.001$, and time step $dt=0.01$. We consider two different types of agent-substrate damping: isotropic damping with mobility coefficients $\mu\equiv\mu_{\parallel}=\mu_{\perp}$ and $\mu=0.02$ (in panels a-d) and fully anisotropic damping with mobility coefficients $\mu_{\perp}=0$ and $\mu_{\parallel}=0.02$ (in panels e-h). The four panels on the left side (a,b,e,f) have a smooth confining boundary while those on the right side (c,d,g,h) have a rough boundary made of inert, unmovable disks. We consider two values for the off-centered rotational distance $R$: large ($R=0.3$) in (a,c,e,g) and small ($R=0.03$) in (b,d,f,h). The temporal dynamics associated with these snapshots are included as Supplementary Movies 1-8. Moreover, the general phase diagrams of this system as a function of $R$ and $D_{\theta}$ in presented in Supplementary Figure 1.
In addition to the three collective states described above, for higher values of noise $D_{\theta}$ and off-centered rotation distance $R$ we observed a standard state of dynamic disorder (where both $P$ and $M$ are very small and agents rotate locally in a disordered way) and a novel state of quenched disorder previously analyzed in [54, 55].
Local and global rotation modes
In order to explore the different rotational modes more in detail, we will examine here the temporal autocorrelation function of the heading direction $\hat{n}_{i}$, which is defined by $C(\tau)=\left\langle\hat{n}_{i}\left(t\right)\cdot\hat{n}_{i}\left(t+\tau\right)\right\rangle_{N,N_{S}}$. Here, $\langle\cdot\rangle_{N,~{}N_{S}}$ is the mean value, averaged over the total number of sampling points $N_{S}=2000$ and over all $N$ agents (see Methods section). We will use this tool to examine in detail the rotational dynamics for the same parameters used in the previous section, but with zero noise ($D_{\theta}=0$), in order to visualize more clearly all persistent oscillations.
@@
Figure 3 plots the orientation autocorrelation functions $C(\tau)$ for the same parameter combinations displayed in Fig. 2. Here, the large $R=0.3$ case is displayed as blue dashed lines, and the small $R=0.03$ case as red solid lines. We begin by noting that the amplitude of the correlations decay rapidly in the rough boundary setups, as shown in panels (b) and (d). This results from the fact that disk-boundary interactions destroy the temporal correlations of agents near the edge. We note, in addition, that the amplitude of the $C(\tau)$ curves with small $R=0.03$ also decay more for systems with anisotropic damping, when compared to their isotropic counterparts. In the smooth boundary systems presented in panels (a) and (c), the angular dynamics appear highly organized, displaying persistent rotations; in the rough boundary system in panels (b) and (d), the temporal orientational correlations rapidly decay to low amplitude. For each of these cases, anisotropic damping produces correlation curves that have lower amplitude and more noise than isotropic damping. All panels show that, for small $R=0.03$, the red curves display a superposition of high and low frequencies. Instead, for large $R=0.3$, the dashed blue lines show only low-frequency oscillations. The orientation autocorrelation functions for a larger system size of $N=1600$ are shown in Supplementary Figure 2.
@@
For a large off-centered distance $R=0.3$, Fig. 4 shows no high frequency peak, so the localized rotation state is fully suppressed, while still displaying the low-frequency maximum that characterizes the milling state. As in the small $R$ case, this maximum is more pronounced in simulations with a smooth boundary, which favor solid body rotation. In contrast to the small $R$ case, the maximum seems to be equally pronounced for isotropic or anisotropic damping, which is likely due to the fact that the motion of an agent with large $R$ is already constrained by external forces to be along its $\hat{n}_{i}$. The Fourier transform of orientation autocorrelation functions for a larger system size of $N=1600$ are shown in Supplementary Figure 3.
Vortex dynamics
@@
The results presented above show that the localized rotation state and the milling state can coexist in some regimes. Moreover, for the parameters used in panels (b) and (d) of Fig. 2, the circular arena divides into two well-structured regions, an outer ring with milling dynamics and a central circular area with localized rotation, with both areas separated by a rotating vortex. We will begin our analysis of this regime by examining in detail the dynamics of this vortex in the anisotropic case (which we verified is equivalent to its dynamics in the isotropic case).
@@
Figure 5 presents the vortex dynamics for the smooth boundary case displayed in Fig. 2(f), with $R=0.03$, $D_{\theta}=0.001$, $N=400$, $l_{0}=1.0$, and $v_{0}=0.002$. After identifying the vortex location, as described in the Methods section, we trace its successive positions every $\delta t=100$ simulation time units ($10^{4}$ simulation time steps). Panel (a) presents the resulting vortex trajectory, showing a single vortex orbiting at an approximately constant rate and fixed distance from the center of the arena. Panel (b) shows that the corresponding angular positions match an orbital speed given by $0.0108$ radians per simulation time unit.
@@
Figure 5: Trajectory and dynamics of orbiting vortex. The position of the orbiting vortex that appears in small $N=400$ systems with a smooth boundary and small off-centered rotation distance $R=0.03$ is displayed over time. a) The trajectory of the vortex is traced in the arena by labeling its position every $\Delta t=100$ computational time units as successive blue dots connected by red arrows. b) Fitting of the cosine function of the angular position of the vortex as a function of time. The blue dots label again the angular position every $\Delta t=100$, while the red curve is a sinusoidal fit given by $\cos{\theta_{v}}=\cos{(0.0108*t+2.03)}$. Other simulation parameters are set to: $\mu_{\parallel}=0.02$ and $\mu_{\perp}=0$ (imposing anisotropic damping), $l_{0}=1.0$, $v_{0}=0.002$, and $D_{\theta}=0.001$.
@@
When we extend our study to the rough boundary case and larger systems, we find that this simple, well structured regime with coexisting regimes separated by a single orbiting vortex becomes unstable. Instead, multiple vortices spontaneously appear and disappear in different parts of the arena. Figure 6 displays the probability density function (PDF) of the radial position all vortices with respect to the center of the arena for $N=400,1600,6400$ agents, keeping the density constant by increasing proportionally the size of the arena. The corresponding scatter plots of all identified vortex positions are displayed in Supplementary Figures 4-7.
@@
We find that only the smooth boundary case with $N=400$ agents present a single vortex in the arena (see Supplemental Figures 4 and 6). For $N=1600$ we find between one and two vortices. For $N=6400$ we always find more than two vortices in the arena, and the number depends on $R$. For $R=0.3$, we find about two vortices in the arena in all cases, whereas for $R=0$, we find close to $5$ vortices in average for isotropic damping and $7$ for anisotropic damping. In general, larger systems, low $R$ values, and anisotropic damping tend to produce more vortices. For larger $R$, the vortices tend to move towards the center of the arena, and for smaller $R$ they form at an approximately fixed distance from the edge.
@@
In the rough boundary case, more vortices tend to nucleate near the boundary. As shown in Supplementary Figures 5 and 7, only the small $N=400$ and large $R=0.3$ cases produce one to two vortices per frame, since here vorticity is mainly driven by the global milling rotation. More vortices appear in the arena for larger $N$ and smaller $R$, reaching an average of close to $13$ and $26$ vortices per frame for $N=6400$ and $R=0$, in cases with isotropic and anisotropic damping, respectively. Here again, more vortices are produced for large $N$ and small $R$. For $R=0.3$ the vortices appear in the bulk of the system but for all larger $R$ values they appear at the edge, right next to the rough boundary.
@@
State transitions as a function of off-centered distance $R$
To complete our analysis of the rotational dynamics, we now consider the transition between different rotational states as a function of the off-centered distance $R$ between the center of rotation and the geometrical center of each agent. We will focus here on the dynamics represented by the largest angular autocorrelation frequency components.
Despite the differences in vortex dynamics described in previous sections, our angular autocorrelation analysis shows that low and high rotational frequencies are also present in larger systems (see Supplementary Figures 3 and 4). Low-frequency angular correlations correspond here to the slow drift of all agents around the center of the arena, whereas the high-frequency ones still reflect localized rotation. For low $R$ values, localized rotation coexists with collective milling. In small $N=400$ systems with smooth boundary, they segregate to produce milling in the outer ring and localized rotation in the inner region, separated by a rotating vortex. In larger systems and for a rough boundary, they overlap without a clear structure.
We examined the presence of high-frequency localized rotation as a function of the off-centered rotation distance $R$ by examining the corresponding peak in the Fourier transform of the temporal angular autocorrelation function (see Supplementary Figures 8 and 9). We found that localized rotation is always present for $R<0.05$, regardless of system size. It is suppressed for higher $R$, where collective milling is favored. We also considered low-frequency milling around the center of the arena. Supplementary Figures 8 and 9 reveal that it is always present and that it displays a relatively constant angular momentum for all $R$, with a similar value for all considered setups (isotropic or anisotropic damping, rough or smooth boundary).
State transitions as a function of density

### v3

@@ Abstract
We study the emerging collective states in a simple mechanical model of a dense group of self-propelled polar disks with off-centered rotation, confined within a circular arena. Each disk presents self-alignment towards the sum of contact forces acting on it, resulting from disk-substrate interactions, while also displaying mutual alignment with neighbors due to having its center of rotation located a distance $R$ behind its centroid, so that central contact forces can also introduce torques. The effect of both alignment mechanisms produces a variety of collective states that combine high-frequency localized circular oscillations with low-frequency milling around the center of the arena, in fluid or solid regimes. We consider cases with small/large $R$ values, isotropic/anisotropic disk-substrate damping, smooth/rough arena boundaries, different densities, and multiple systems sizes, showing that the emergent collective states that we identify are robust, generic, and potentially observable in real-world natural or artificial systems.
Copyright attribution to authors.
This work is a submission to SciPost Physics.
License information to appear upon publication.
Publication information to appear upon publication.
Received Date
Accepted Date
Published Date
1 Introduction
@@
A natural way to combine self-alignment and mutual alignment effects in a mechanical model is to consider off-centered attraction-repulsion forces that are not radial with respect to the center of rotation. The inter-agent forces then introduce torques that can result in mutual alignment, while simultaneously producing self-alignment. This approach was implemented in a minimal model of active solid introduced in [54], where a system of self-propelled agents connected by springs attached to lever arms was shown to display two types of order-disorder transition, one driven by the focusing of the self-propulsion energy into low elastic modes (produced by self-alignment) and the other by the decentralized heading consensus (resulting from mutual alignment). A similar approach was carried out to show that both types of self-organizing dynamics are also present in a minimal model describing densely packed self-propelled disks with linear repulsion only, each with an off-centered center of rotation, behind its geometrical center [55].
@@
Our work shows that dense systems of self-propelled agents with off-centered, nonaxisymmetric rotational dynamics can produce, both, self-alignment and mutual alignment processes, demonstrating the coexistence of two forms of self-organization in active systems. We expect experimental systems to be capable of displaying both types of self-organizing dynamics in real-world setups (since off-centered and nonaxisymmetric rotation will naturally occur if the interactions between neighbors and with the substrate are not perfectly axisymmetric), potentially resulting in a rich combination of emergent states.
2 Model Setup
2.1 Dynamics of active polar disks with off-centered rotation
We describe a dense system of active elastic agents confined in a circular arena, to carry out simulations as detailed in Appendix A. Each agent is a soft disk $i$ with position $\vec{r}_{i}$, defined by its center of rotation that self-propels with preferred speed $v_{0}$ along the heading direction indicated by its orientation unit vector $\hat{n}_{i}=(\cos(\theta_{i}),\sin(\theta_{i}))$, as shown in Fig. 1a. All disks have the same diameter $l_{0}$ and a repulsive force appears between disks $i$ and $j$ when they overlap, given by $\vec{f}_{ij}=(k/l_{0})(|\vec{l}_{ij}|-l_{0})\vec{l}_{ij}/|\vec{l}_{ij}|$ if $|\vec{l}_{ij}|<l_{0}$, and by $\vec{f}_{ij}=0$ otherwise, where $\vec{l}_{ij}=\vec{l}_{j}-\vec{l}_{i}$ with $\vec{l}_{i}=\vec{r}_{i}+R\hat{n}_{i}$. The total repulsive forces exerted by other agents on disk $i$ will thus be $\vec{F}_{i}={\textstyle\sum_{j\epsilon S_{i}}}\vec{f}_{ij}$, where $S_{i}$ includes the indexes of all the neighbors of disk $i$. The combined effect of the interaction forces and self-propulsion leads to the following overdamped equation of motion for the position of each agent [55, 4]
@@ 2.1 Dynamics of active polar disks with off-centered rotation
The disk-substrate interaction can lead to self-alignment, where the heading tends to align towards the sum of total forces on the disk. While self-aligning torques may originate from various mechanisms, as detailed in [4], we consider here that they have the form $(\hat{n}_{i}\times\vec{F}_{i})\times\hat{n}_{i}=(\hat{n}_{i}\times\vec{F}_{i})\cdot\hat{z}=(\vec{F}_{i}\cdot\hat{n}_{i}^{\perp})\hat{n}_{i}^{\perp}=(\mathbb{I}-\hat{n}_{i}\hat{n}^{T}_{i})\vec{F}_{i}$, where $\hat{z}$ is the unit vector normal to the plane and $\beta$ controls the self-aligning strength. Note that these torques will depend on the details of the disk-substrate interaction. On the other hand, the position and center of rotation $\vec{r}_{i}$ of each disk $i$ is located a distance $R$ behind its geometrical center (with respect to its heading direction), at $\vec{l}_{i}=\vec{r}_{i}+R\hat{n}_{i}$, which introduces mutual alignment due to the torques on the arm of length $R$ between the geometrical center (where the central forces are acting) and the center of rotation. By combining these torques with the effects of noise, we find a dynamical equation for the heading direction $\hat{n}_{i}$, given by
@@
where $\eta_{i}$ is a random variable that introduces Gaussian white noise with zero mean and variance $\left\langle\eta_{i}(t)\eta_{j}(t^{\prime})\right\rangle=\delta_{ij}\delta(t-t^{\prime})$. Here, $D_{\theta}$ sets the rotational noise strength and corresponds to the rotational diffusion coefficient.
Figure 1: Active polar disks with linear elastic repulsion model and arena. a Schematic diagram of two interacting self-propelled disks, $i$ and $j$, oriented along $\hat{n}_{i}$ and $\hat{n}_{j}$, respectively. Each disk has diameter $l_{0}$ and its center of rotation is labeled by a red point, located a distance $R$ behind its geometrical center. Linear repulsive forces, proportional to $|\vec{l}_{ij}|-l_{0}$, act on each centroid, resulting in forces and torques on $\vec{r}_{i}$ and $\vec{r}_{j}$. b Snapshot of a relatively ordered densely packed group of active polar disks in circular confinement, inside a smooth circular barrier introduced as a radial linear outer potential. c Snapshot of a relatively disordered densely packed group of active polar disks confined by a ring of fixed disks with the same repulsion potential as the active disks, representing a rough boundary. Each arrow corresponds to an active disk and its self-propulsion direction, colored by angle.
2.2 Simulation arena and parameters
We study the collective regimes that result from integrating Eqs. (1) and (2) for a dense group of disks confined in a circular arena under different conditions. The simulation arena is defined by a smooth or rough circular confinement barrier. In the smooth circular confinement case, the outer ring boundary exerts a repulsive radial linear force inwards, perpendicular to its surface, as shown in Fig. 1b. In the rough circular confinement case, the boundary is composed of fixed disks with a repulsive potential equivalent to that of the active disks in the simulation, as shown in Fig. 1c.
We consider homogeneous soft disks, all having the same diameter $l_{0}=1$ and self-propulsion speed of $v_{0}=0.002$, unless otherwise noted. The force alignment strength coefficient is set to $\beta=1.2$ and the simulation time step is $dt=0.01$ in all simulations. All simulations are carried out starting from random initial positions and orientations.
To characterize different scenarios, two values of the off-centered rotation distance $R$ are chosen: $R=0.03$ for weak mutual alignment $R=0.3$ and for strong mutual alignment. The $R$-dependent transition in vortex dynamics is further explored by varying $R$ from 00 to $0.3$. We consider two limiting cases of mobility: a fully isotropic case with $\mu\equiv\mu_{\parallel}=\mu_{\perp}$ and $\mu=0.02$, where the disk-substrate damping is uniform across all directions, and a fully anisotropic damping case with $\mu_{\perp}=0$ and $\mu_{\parallel}=0.02$, where the agents can only move forward or backward (as in the case of wheeled vehicles).
Most analyses focus on systems with $N=400$ agents, while additional simulations with $N=1600$ and $N=6400$ agents are conducted to examine how the emergent states evolve in larger systems. The radius of the circular arena is set to $R_{c}=10.5$ for $N=400$ soft polar disks, leading to a packing fraction $\phi=Nl_{0}^{2}/4R_{c}^{2}=0.907$, where $l_{0}=1$ is the disk diameter. We vary the radius of the circular confinement, $R_{c}$, for different total numbers of active polar disks, $N$, to maintain constant packing fraction, $\phi=Nl_{0}^{2}/4R_{c}^{2}$, as shown in Fig. 6. For $N=400,1600,6400$, we set $R_{c}=10.5,21,42$ with $l_{0}=1$, all corresponding to $\phi=0.907$.
A detailed definition of all the simulation variables, along with their specific values in our simulations, are provided in Table 1. The parameter lists used to generate the majority of the simulation results presented in this article. Any difference with respect to these values is explicitly noted in the figure captions.
3 Results
3.1 Overview of emergent states
We begin by providing an overview of the different regimes observed for agents with isotropic or anisotropic damping and small or large $R$, confined by smooth or rough boundaries. All simulations were carried out as detailed in Appendix A.
We characterize the emerging collective states using two order parameters. The first is the commonly used Polarization ($P$). It shows the degree of alignment and is defined by
$P=\frac{1}{N}\langle\|\sum_{i=1}^{N}{\hat{n}_{i}}\|\rangle_{t}~{},$ (3)
where $\left\langle\cdot\right\rangle_{t}$ represents the time average after reaching steady state. The second one measures the degree of milling in the system, and is defined as the milling order parameter
Here, $\hat{r}_{ig}=\left(\vec{r}_{i}-\vec{r}_{g}\right)/\left\|\vec{r}_{i}-\vec{r}_{g}\right\|$ points to the position of agent $i$ with respect to the center of the group $\vec{r}_{g}$. The order parameters $P$ and $M$ range from 00 to $1$, where a disordered state will display low values for both, a flocking state will have high $P$ and low $M$, and a milling state will have low $P$ and high $M$.
Fig. 2 illustrates the collective states observed in simulations of 400 disks with diameter $l_{0}=1$, self-propulsion speed $v_{0}=0.002$, alignment strength $\beta=1.2$, low noise $D_{\theta}=0.001$, and time step $dt=0.01$. Two types of agent-substrate damping are considered: isotropic damping with mobility coefficients $\mu\equiv\mu_{\parallel}=\mu_{\perp}$ and $\mu=0.02$ (in panels a-d) and fully anisotropic damping with $\mu_{\perp}=0$ and $\mu_{\parallel}=0.02$ (in panels e-h). The four panels on the left side (a,b,e,f) correspond to smooth confinement, while the right panels (c,d,g,h) represent rough boundaries composed of inert, immovable disks. Two values for the off-centered rotational distance $R$ are explored: large ($R=0.3$) in (a,c,e,g) and small ($R=0.03$) in (b,d,f,h). The corresponding temporal dynamics are available as Supplementary Movies 1-8.
In addition to the three collective states described above, for higher values of noise $D_{\theta}$ and off-centered rotation distance $R$ we observed a standard state of dynamic disorder (where both $P$ and $M$ are very small and agents rotate locally in a disordered way) and a novel state of quenched disorder previously analyzed in [54, 55], both of which are shown in the phase diagrams displayed in Appendix B.
@@ 3.1 Overview of emergent states
3.2 Local and global rotation modes
In order to explore the different rotational modes more in detail, we compute the orientation temporal autocorrelation function $C(\tau)$ to evaluate how the particles’ $\hat{n}_{i}(t)$ orientations change over time and identify their rotational frequencies. We define this function for $N$ particles using the expression
We will use this tool to examine in detail the rotational dynamics for the same parameters used in the previous section, but with zero noise ($D_{\theta}=0$), in order to visualize more clearly all persistent oscillations.
@@ 3.2 Local and global rotation modes
Figure 3 plots the orientation autocorrelation functions $C(\tau)$ for the same parameter combinations displayed in Fig. 2. Here, the large $R=0.3$ case is displayed as blue dashed lines, and the small $R=0.03$ case as red solid lines. It is notable that the amplitude of correlations decays rapidly in rough boundary setups (panels (b) and (d)), as disk-boundary interactions disrupt the temporal correlations of agents near the edge. We note, in addition, that the amplitude of the $C(\tau)$ curves in the small $R=0.03$ case also decay more rapidly for systems with anisotropic damping, when compared to their isotropic counterparts (i.e., the amplitude of the solid red curves becomes smaller more rapidly in the anisotropic case). In the smooth boundary setups (panels (a) and (c)), angular dynamics are highly organized, displaying persistent rotations. Conversely, in rough boundary setups (panels (b) and (d)), temporal orientational correlations rapidly decay to low amplitude. Across all cases, anisotropic damping produces lower-amplitude correlation curves than isotropic damping. Notably, for small $R=0.03$, red curves exhibit a superposition of high- and low-frequency oscillations, whereas large $R=0.3$ cases (blue dashed lines) display only low-frequency oscillations. The orientation autocorrelation functions for a larger system size of $N=1600$ are shown in Fig. 12.
@@
For a large off-centered distance $R=0.3$, Fig. 4 shows no high frequency peak, so the localized rotation state is fully suppressed, while still displaying the low-frequency maximum that characterizes the milling state. As in the small $R$ case, this maximum is more pronounced in simulations with a smooth boundary, which favor solid body rotation. In contrast to the small $R$ case, the maximum seems to be equally pronounced for isotropic or anisotropic damping, which is likely due to the fact that the motion of an agent with large $R$ is already constrained by external forces to be along its $\hat{n}_{i}$. The Fourier transform of orientation autocorrelation functions for a larger system size of $N=1600$ are shown in Fig. 13. Finally, we found that the orientation autocorrelation function has the same frequency components in different regions in the bulk, so the rotational dynamics of the disks are essentially homogeneous throughout the arena until the boundary region is reached.
3.3 Vortex dynamics
@@
The results presented above show that the localized rotation state and the milling state can coexist in some regimes. Moreover, for the parameters used in panels (b) and (d) of Fig. 2, the circular arena divides into two distinct regions: an outer ring exhibiting milling dynamics and a central circular area characterized by localized rotation, separated by a rotating vortex. We begin our analysis of this regime by examining the dynamics of this vortex in the anisotropic case, which is equivalent to its behavior in the isotropic case.
@@
Figure 5 presents the vortex dynamics for the smooth boundary case displayed in Fig. 2f, with $R=0.03$, $D_{\theta}=0.001$, $N=400$, $l_{0}=1.0$, and $v_{0}=0.002$. After identifying the vortex location, as detailed in Appendix C, its successive positions are traced every $\delta t=100$ simulation time units ($10^{4}$ simulation time steps). Panel (a) depicts the vortex trajectory, showing a single vortex orbiting at an approximately constant rate and fixed distance from the center of the arena. Panel (b) shows that the corresponding angular positions match an orbital speed given by $0.0108$ radians per simulation time unit.
@@
Figure 5: Trajectory and dynamics of orbiting vortex. The position of the orbiting vortex that appears in small $N=400$ systems with a smooth boundary and small off-centered rotation distance $R=0.03$ is displayed over time. a The trajectory of the vortex is traced in the arena by labeling its position every $\Delta t=100$ computational time units as successive blue dots connected by red arrows. b Fitting of the cosine function of the angular position of the vortex as a function of time. The blue dots label again the angular position every $\Delta t=100$, while the red curve is a sinusoidal fit given by $\cos{\theta_{v}}=\cos{(0.0108*t+2.03)}$. Other simulation parameters are set to: $\mu_{\parallel}=0.02$ and $\mu_{\perp}=0$ (imposing anisotropic damping), $l_{0}=1.0$, $v_{0}=0.002$, and $D_{\theta}=0.001$.
@@
When we extend our study to the rough boundary case and larger systems, we find that this simple, well structured regime with coexisting regimes separated by a single orbiting vortex becomes unstable. Instead, multiple vortices spontaneously appear and disappear in different parts of the arena. Fig. 6 displays the probability density function (PDF) of the radial position all vortices with respect to the center of the arena for $N=400,1600,6400$ agents, keeping the density constant by increasing proportionally the size of the arena. The corresponding scatter plots of all identified vortex positions are displayed in Figs. 1619.
@@
We find that only the smooth boundary case with $N=400$ agents present a single vortex in the arena (see Figs. 16 and 18). For $N=1600$ we find between one and two vortices. For $N=6400$ we always find more than two vortices in the arena, and the number depends on $R$. For $R=0.3$, we find about two vortices in the arena in all cases, whereas for $R=0$, we find close to $5$ vortices on average for isotropic damping and $7$ for anisotropic damping. In general, larger systems, low $R$ values, and anisotropic damping tend to produce more vortices. For larger $R$, the vortices tend to move towards the center of the arena, and for smaller $R$ they form at an approximately fixed distance from the edge.
@@
In the rough boundary case, more vortices tend to nucleate near the boundary. As shown in Fig. 17 and  19, only the small $N=400$ and large $R=0.3$ cases produce one to two vortices per frame, since here vorticity is mainly driven by the global milling rotation. More vortices appear in the arena for larger $N$ and smaller $R$, reaching an average of close to $13$ and $26$ vortices per frame for $N=6400$ and $R=0$, in cases with isotropic and anisotropic damping, respectively. Here again, more vortices are produced for large $N$ and small $R$. For $R=0.3$ the vortices appear in the bulk of the system but for all larger $R$ values they appear at the edge, right next to the rough boundary.
@@
3.4 State transitions as a function of off-centered distance $R$
To complete our analysis of the rotational dynamics, we now consider the transition between different rotational states as a function of the off-centered distance $R$ between the center of rotation and the geometrical center of each agent. The full phase diagrams in $D_{\theta}-R$ plane are shown in Appendix B (Fig. 9). We will focus here on the dynamics represented by the largest angular autocorrelation frequency components.
Despite the differences in vortex dynamics described in previous sections, our angular autocorrelation analysis shows that low and high rotational frequencies are also present in larger systems (see Figs. 12 and 13). Low frequency angular correlations correspond here to the slow drift of all agents around the center of the arena, whereas the high frequency ones still reflect localized rotation. For low $R$ values, localized rotation coexists with collective milling. In small $N=400$ systems with smooth boundary, they segregate to produce milling in the outer ring and localized rotation in the inner region, separated by a rotating vortex. In larger systems and for a rough boundary, they overlap without a clear structure.
As shown in Figs. 4 and 13 of the paper, low- and high-frequency peaks in the orientation temporal autocorrelation functions can be identified for different values of $R$.
By implementing a peak detection algorithm on the Fourier transformed curves of the orientation autocorrelation, we can display the frequency and amplitude of each peak as a function of $R$, the parameter that controls the off-centered rotation distance, and thus the level of mutual alignment compared to self-alignment.
The Fig. 14 display the frequency and amplitude of the main low- and high-frequency peaks as a function of $R$ for different combinations of isotropic (I) or anisotropic (A) agent-substrate interactions and a smooth (S) or rough (R) confining boundary, as indicated by the column titles. This figure presents the small system case with $N=400$, whereas Fig. 15 displays large system case with $N=1600$, keeping the same mean density or packing fraction.
We found that localized rotation is always present for $R<0.05$, regardless of system size. It is suppressed for higher $R$, where collective milling is favored. We also considered low-frequency milling around the center of the arena. Fig. 14 and 15 reveal that it is always present and that it displays a relatively constant angular momentum for all $R$, with a similar value for all considered setups (isotropic or anisotropic damping, rough or smooth boundary).
3.5 State transitions as a function of density